# Large Scale Transfer Learning for Tabular Data via Language Modeling

### v1

@@ G.2 Impact of Contamination
Figure 11: 11(a): Results curves on the subset of tasks identified as potentially contaminated according to our fuzzy decontamination procedure. 11(b): Results curves on the subset of tasks not identified as potentially contaminated according to our fuzzy decontamination procedure.
@@ I.2 Intended Use
Out-of-scope use cases: Commercial use, use of the model to attenmpt to identify, harm, or violate the privact of individuals represented in the training data, and any other behavior that violates the Meta Llama 3 license is out of scope.
@@ I.5 Evaluation Data
Datasets: We use a suite of five previously-proposed tabular benchmarks, conprising a total of over 300 tables. Our evaluation datasets are described in detail in Sections 5.2 and D.1.
@@ I.9 Caveats and Recommendations
The dataset consists of 3.1M tables where each table has many rows. The total number of rows across tables is 1.6B.

### v2

@@ F.5 Ablation Study: Data Filtering
F.6 Ablation Study: Base Language Model
In this section, we evaluate the improvements of TabuLa-8B due to improvements in the underlying base language model. In order to do so, we train variants of TabuLa-8B which are identical to the final version, except we use Llama 1 and Llama 2 as the base language models. This allows us to investigate the improvements in TabuLa-8B as the underlying language model improves. We compare the Llama 1 and Llama 2 variants to a compute-matched Llama 3 variant (we use only 10% of the compute relative to our final model, as in our other ablation studies, but also compare to the 100%-compute TabuLa-8B for reference). We note that, due to the smaller context sizes of the Llama 1 and Llama 2 models, we ensure the comparisons are example-matched; that is, we train Llama 2 (context size 4096) for 2x the update steps relative to Llama 3 (context size 8192), and train Llama 1 (context size 2048) for 4x the update steps. This ensures that all models see roughly the same number of tokens or examples during fine-tuning.
The results of this study are shown in Figure 16. We only report the results up to 8 shots to allow for fair comparisons across all models, as Llama 1’s context size of 2048 is 4x smaller than TabuLa-8B and can only fit up to 8 shots for many tasks. Figure 16 shows the clear improvement from better base models: as the underlying base models improve (Llama 1 $\rightarrow$ Llama 2 $\rightarrow$ Llama 3), the fine-tuned model also improves. These results provide hope that further improvements in language modeling could also lead to gains in tabular models.
Figure 16: Results of base model ablation study.
@@ G.2 Impact of Contamination
Figure 17: 17(a): Results curves on the subset of tasks identified as potentially contaminated according to our fuzzy decontamination procedure. 17(b): Results curves on the subset of tasks not identified as potentially contaminated according to our fuzzy decontamination procedure.
@@ I.2 Intended Use
Out-of-scope use cases: Commercial use, use of the model to attempt to identify, harm, or violate the privacy of individuals represented in the training data, and any other behavior that violates the Meta Llama 3 license is out of scope.
@@ I.5 Evaluation Data
Datasets: We use a suite of five previously-proposed tabular benchmarks, comprising a total of 329  tables. Our evaluation datasets are described in detail in Sections 5.2 and D.1.
@@ I.9 Caveats and Recommendations
The dataset consists of 4M tables where each table has many rows. The total number of rows across tables is 2.1B.